# When is Truthfully Allocating Chores no Harder than Goods?

### v1

@@ 4.1 Characterization
A swap-dictatorial mechanism $\mathcal{M}$ for divisible goods (resp. chores) is defined by a (potentially infinite) set of bundles $D\subseteq[0,1]^{m}$ such that for every valuation profile $\mathbf{v}$ (resp. cost profile $\boldsymbol{c}$), $\mathcal{M}_{1}(\mathbf{v})=(x_{1}+\mathbf{1}_{O}-x_{2})/2$ (resp. $\mathcal{M}_{1}(\boldsymbol{c})=(x_{1}+\mathbf{1}_{O}-x_{2})/2$), where $x_{1}\in\mathop{\arg\max}_{x\in D}v_{1}(x)$ (resp. $x_{1}\in\mathop{\arg\min}_{x\in D}c_{1}(x)$) and $x_{2}\in\mathop{\arg\max}_{x\in D}v_{2}(x)$ (resp. $x_{2}\in\mathop{\arg\min}_{x\in D}c_{2}(x)$) with ties broken arbitrarily.
@@
A mechanism $\mathcal{M}$ for divisible items is smooth if, for every $i\in N$, $\mathcal{M}_{i}(\mathbf{v})$ is twice continuously differentiable with respect to every utility function $v_{j}$, $j\in N$ [^5] Here, we encode a utility function $v_{j}$ by the vector $(v_{j}(o_{1}),\ldots,v_{j}(o_{m}))\in[0,1]^{m}$, and we say that a condition holds with respect to a utility function if it holds with respect to the corresponding vector in $[0,1]^{m}$. The same definition holds for cost profiles.
We first recall the characterization of truthful mechanisms for goods by [FWVP23].
Theorem 4.5 ([FWVP23]).
For divisible items with two agents, an anonymous and smooth mechanism is truthful for goods iff it is swap-dictatorial for goods.
@@
For divisible items with two agents, an anonymous and smooth mechanism is truthful for chores iff it is swap-dictatorial for chores.
@@ 4.2 Efficiency Ratio
For divisible items with two agents, if a mechanism $\mathcal{M}$ is smooth and truthful for chores, then $\mathcal{M}$ has an efficiency ratio of 00 for chores.
It is worth of mentioning that [BHS20] show that in cake-cutting with two agents, the efficiency ratio of every EF and truthful mechanism for chores has an efficiency ratio of 00 for chores. However, Theorem 4.7 is arguably stronger than their result since (1) agents have less room to manipulate for homogeneous divisible items than in cake-cutting, and (2) Theorem 4.7 holds even for non-EF mechanisms.
@@
For divisible items with two agents, given a truthful mechanism with an efficiency ratio of $\delta$ for chores, there exists a corresponding anonymous, item-symmetric, and truthful mechanism with an efficiency ratio of at least $\delta$ for chores.
@@
Assume that items are divisible. Let $D$ be the dictator’s choice space of a swap-dictatorial mechanism (for either goods or chores) that is anonymous and item-symmetric. If a bundle $x=(x(o_{1}),\ldots,x(o_{m}))\in D$, then $(x(\sigma(o_{1})),\ldots,x(\sigma(o_{m})))\in D$ for every $\sigma\in\mathsf{Sym}(O)$.
@@ 5 Indivisible Items with Bi-Valued Preferences
We first recall the truthful, PO, and EF mechanism of [BT24] for divisible goods with bi-valued utility functions, denoted as $\mathcal{M}^{g}$. Given a bi-valued utility profile $\mathbf{v}$ as input, where $v_{i}(o)\in\{p,q\}$ for all $i\in N$ and $o\in O$ with $p>q>0$, we assume that for each agent $i$, there exists an item $o$ such that $v_{i}(o)=p$, which is without loss of generality since if $v_{i}(o)=q$ for every item $o$, then we can equivalently treat the utility function of agent $i$ to satisfy $v_{i}(o)=p$ for every item $o$. Let $\mathbf{x}^{\prime}$ be an arbitrary allocation that maximizes
@@
Notice that $x^{\prime}_{i}(o)>0$ implies $v_{i}(o)=p$ for all $i\in N$ and $o\in O$ by the optimality of $\mathbf{x}^{\prime}$, and $f_{\mathbf{v}}(\mathbf{x}^{\prime})>0$ since there exists at least one item $o$ for each agent $i$ such that $v_{i}(o)=p$. Let $L=m/n$, and we prove a structural property of $\mathbf{x}^{\prime}$ in the following lemma.
@@
For every $o\in O$, let $N_{o}=\{i\in N\mid x_{i}^{\prime}(o)>0\}$ be the set of agents receiving some fraction of $o$ in $\mathbf{x}^{\prime}$. Then, either $|x_{i}^{\prime}|>L$ for every $i\in N_{o}$, or $|x_{i}^{\prime}|\leq L$ for every $i\in N_{o}$. Moreover, if $|x_{i}^{\prime}|>L$ for some $i\in N_{o}$, then $v_{j}(o)=q$ for every $j\in N$ such that $|x_{j}^{\prime}|\leq L$.
@@
Now, we construct a price vector $\mathbf{r}$ for the allocation $\mathbf{x}=\mathcal{M}^{c}(\boldsymbol{c})$ and assign a budget to each agent so that $(\mathbf{x},\mathbf{r})$ constitutes a market equilibrium. Recall that $\mathbf{v}$ is the bi-valued utility profile with $v_{i}(o)=(p+q)-c_{i}(o)$ for all $i\in N$ and $o\in O$, $\mathbf{x}^{\prime}$ is the allocation that maximizes $f_{\mathbf{v}}(\cdot)$ defined by (3), and $\mathbf{x}^{\prime\prime}$ is the allocation obtained from $\mathbf{x}^{\prime}$ by truncating the bundles $x_{i}^{\prime}$ with $|x_{i}^{\prime}|>L$. Let $Z=\{i\in N\mid|x_{i}^{\prime}|>L\}$ denote the set of agents whose bundles are truncated. For each item $o$, we know by the first statement in Lemma 5.4 that in $\mathbf{x}^{\prime}$, either $o$ is fully allocated to agents in $N\setminus Z$, in which case we set the price of $r(o)=q$, or $o$ is fully allocated to agents in $Z$, in which case we set $r(o)=p$. Moreover, each agent $i$’s budget $b_{i}$ is defined as the total price of $x_{i}$, i.e.,
@@
Let $\boldsymbol{c}$ be an additive cost profile and $\mathbf{x}$ be a fractional allocation corresponding to an outcome of PSc with respect to $\boldsymbol{c}$. Let $c_{\max}=\max_{i\in N,o\in O}c_{i}(o)>0$, and we define an additive utility profile $\mathbf{v}$ with $v_{i}(o)=c_{\max}-c_{i}(o)$ for all $i\in N$ and $o\in O$. Since $v_{i}(o)\geq v_{i}(o^{\prime})$ iff $c_{i}(o)\leq c_{i}(o^{\prime})$ for all $i\in N$ and $o,o^{\prime}\in O$, it suffices to find a random integral allocation $\widehat{A}$ such that $\Pr[o\in\widehat{A}_{i}]=x_{i}(o)$ for all $i\in N$ and $o\in O$. Moreover, for each integral allocation $A$ in the support of $\widehat{A}$, it satisfies that $|A_{i}|\leq|A_{j}|+1$ for all $i,j\in N$, and we can label the items in each bundle $A_{i}$ as $o_{i}^{1},o_{i}^{2},\ldots,o_{i}^{|A_{i}|}$ such that
@@ Acknowledgments
The research of Biaoshuai Tao was supported by the National Natural Science Foundation of China (Grant No. 62472271).

### v2

@@ 4.1 Characterization
A swap-dictatorial mechanism $\mathcal{M}$ for homogeneous divisible goods (resp. chores) is defined by a (potentially infinite) set of bundles $D\subseteq[0,1]^{m}$ such that for every valuation profile $\mathbf{v}$ (resp. cost profile $\boldsymbol{c}$), $\mathcal{M}_{1}(\mathbf{v})=(x_{1}+\mathbf{1}_{O}-x_{2})/2$ (resp. $\mathcal{M}_{1}(\boldsymbol{c})=(x_{1}+\mathbf{1}_{O}-x_{2})/2$), where $x_{1}\in\mathop{\arg\max}_{x\in D}v_{1}(x)$ (resp. $x_{1}\in\mathop{\arg\min}_{x\in D}c_{1}(x)$) and $x_{2}\in\mathop{\arg\max}_{x\in D}v_{2}(x)$ (resp. $x_{2}\in\mathop{\arg\min}_{x\in D}c_{2}(x)$) with ties broken arbitrarily.
@@
A mechanism $\mathcal{M}$ for homogeneous divisible items is smooth if, for every $i\in N$, $\mathcal{M}_{i}(\mathbf{v})$ is twice continuously differentiable with respect to every utility function $v_{j}$, $j\in N$ [^5] Here, we encode a utility function $v_{j}$ by the vector $(v_{j}(o_{1}),\ldots,v_{j}(o_{m}))\in[0,1]^{m}$, and we say that a condition holds with respect to a utility function if it holds with respect to the corresponding vector in $[0,1]^{m}$. The same definition holds for cost profiles.
With the above definitions, we are ready to present the characterization of truthful mechanisms for goods by [FWVP23].
Theorem 4.5 (Theorem 6 in [FWVP23]).
For homogeneous divisible items with two agents, an anonymous and smooth mechanism is truthful for goods iff it is swap-dictatorial for goods.
@@
For homogeneous divisible items with two agents, an anonymous and smooth mechanism is truthful for chores iff it is swap-dictatorial for chores.
@@ 4.2 Efficiency Ratio
For homogeneous divisible items with two agents, if a mechanism $\mathcal{M}$ is smooth and truthful for chores, then $\mathcal{M}$ has an efficiency ratio of 00 for chores.
It is worth mentioning that [BHS20] show that in cake-cutting with two agents, the efficiency ratio of every EF and truthful mechanism for chores has an efficiency ratio of 00 for chores. However, Theorem 4.7 is arguably stronger than their result since (1) agents have less room to manipulate for homogeneous divisible items than in cake-cutting, and (2) Theorem 4.7 holds even for non-EF mechanisms.
@@
For homogeneous divisible items with two agents, given a truthful mechanism with an efficiency ratio of $\delta$ for chores, there exists a corresponding anonymous, item-symmetric, and truthful mechanism with an efficiency ratio of at least $\delta$ for chores.
@@
Assume that items are divisible and homogeneous. Let $D$ be the dictator’s choice space of a swap-dictatorial mechanism (for either goods or chores) that is anonymous and item-symmetric. If a bundle $x=(x(o_{1}),\ldots,x(o_{m}))\in D$, then $(x(\sigma(o_{1})),\ldots,x(\sigma(o_{m})))\in D$ for every $\sigma\in\mathsf{Sym}(O)$.
@@ 5 Indivisible Items with Bi-Valued Preferences
We first recall the truthful, PO, and EF mechanism of [BT24] for divisible goods with bi-valued utility functions, denoted as $\mathcal{M}^{g}$. Given a bi-valued utility profile $\mathbf{v}$ as input, where $v_{i}(o)\in\{p,q\}$ for all $i\in N$ and $o\in O$ with $p>q>0$, we assume that for each agent $i$, there exists an item $o$ such that $v_{i}(o)=p$, which is without loss of generality since if $v_{i}(o)=q$ for every item $o$, then we can equivalently treat the utility function of agent $i$ to satisfy $v_{i}(o)=p$ for every item $o$. Let $\mathbf{x}^{\prime}$ be an arbitrary (possibly partial) allocation that maximizes
@@
In particular, let $H_{\mathbf{v}}=\{o\in O\mid v_{i}(o)=q\text{ for every }i\in N\}$, and here we do not allocate the items in $H_{\mathbf{v}}$, i.e., $x_{i}^{\prime}(o)=0$ for all $i\in N$ and $o\in H_{\mathbf{v}}$. Notice that $x^{\prime}_{i}(o)>0$ implies $v_{i}(o)=p$ for all $i\in N$ and $o\in O$ by the optimality of $\mathbf{x}^{\prime}$, and $f_{\mathbf{v}}(\mathbf{x}^{\prime})>0$ since there exists at least one item $o$ for each agent $i$ such that $v_{i}(o)=p$. Let $L=m/n$, and we prove a structural property of $\mathbf{x}^{\prime}$ in the following lemma, which will be useful in the proof of Lemma 5.6 to construct a market equilibrium.
@@
For every $o\in O\setminus H_{\mathbf{v}}$, let $N_{o}=\{i\in N\mid x_{i}^{\prime}(o)>0\}$ be the set of agents receiving some fraction of $o$ in $\mathbf{x}^{\prime}$. Then, either $|x_{i}^{\prime}|>L$ for every $i\in N_{o}$, or $|x_{i}^{\prime}|\leq L$ for every $i\in N_{o}$. Moreover, if $|x_{i}^{\prime}|>L$ for some $i\in N_{o}$, then $v_{j}(o)=q$ for every $j\in N$ such that $|x_{j}^{\prime}|\leq L$.
@@
Now, we construct a price vector $\mathbf{r}$ for the allocation $\mathbf{x}=\mathcal{M}^{c}(\boldsymbol{c})$ and assign a budget to each agent so that $(\mathbf{x},\mathbf{r})$ constitutes a market equilibrium. Recall that $\mathbf{v}$ is the bi-valued utility profile with $v_{i}(o)=(p+q)-c_{i}(o)$ for all $i\in N$ and $o\in O$, $\mathbf{x}^{\prime}$ is the allocation that maximizes $f_{\mathbf{v}}(\cdot)$ defined by (3) with items in $H_{\mathbf{v}}$ unallocated, and $\mathbf{x}^{\prime\prime}$ is the allocation obtained from $\mathbf{x}^{\prime}$ by truncating the bundles $x_{i}^{\prime}$ with $|x_{i}^{\prime}|>L$. Let $Z=\{i\in N\mid|x_{i}^{\prime}|>L\}$ denote the set of agents whose bundles are truncated. For each item $o\in O\setminus H_{\mathbf{v}}$, we know by the first statement in Lemma 5.4 that in $\mathbf{x}^{\prime}$, either $o$ is fully allocated to agents in $N\setminus Z$, in which case we set the price $r(o)=q$, or $o$ is fully allocated to agents in $Z$, in which case we set $r(o)=p$. Moreover, for each item $o\in H_{\mathbf{v}}$, we set $r(o)=p$. Finally, each agent $i$’s budget $b_{i}$ is defined as the total price of $x_{i}$, i.e.,
@@
Let $\boldsymbol{c}$ be an additive cost profile and $\mathbf{x}$ be a fractional allocation corresponding to an outcome of PSc with respect to $\boldsymbol{c}$. Let $c_{\max}=\max_{i\in N,o\in O}c_{i}(o)>0$, and we define an additive utility profile $\mathbf{v}$ with $v_{i}(o)=c_{\max}-c_{i}(o)$ for all $i\in N$ and $o\in O$. Since $v_{i}(o)\geq v_{i}(o^{\prime})$ iff $c_{i}(o)\leq c_{i}(o^{\prime})$ for all $i\in N$ and $o,o^{\prime}\in O$, it suffices to find a random integral allocation $\widehat{A}$ such that $\Pr[o\in\widehat{A}_{i}]=x_{i}(o)$ for all $i\in N$ and $o\in O$. Moreover, we will show that the given $\widehat{A}$ also satisfies that for each integral allocation $A$ in the support of $\widehat{A}$, $|A_{i}|\leq|A_{j}|+1$ for all $i,j\in N$, and we can label the items in each bundle $A_{i}$ as $o_{i}^{1},o_{i}^{2},\ldots,o_{i}^{|A_{i}|}$ such that
@@ Acknowledgments
We would like to thank the anonymous reviewers for their detailed feedback and comments that helped significantly improve the exposition of this paper. Bo Li is partly funded by the Hong Kong SAR Research Grants Council (No. PolyU 15224823) and the Department of Science and Technology of Guangdong Province (No. 2023A1515010592). The research of Biaoshuai Tao is supported by the National Natural Science Foundation of China (No. 62472271) and the Key Laboratory of Interdisciplinary Research of Computation and Economics (Shanghai University of Finance and Economics), Ministry of Education. Xiaowei Wu is funded by the Science and Technology Development Fund (FDCT), Macau SAR (file no. 0147/2024/RIA2, 0014/2022/AFJ, 0085/2022/A, and 001/2024/SKL).